# url]https://www.bangalab.org

### v1

@@ Abstract
Living organisms exhibit remarkable adaptations across all scales, from molecules to ecosystems. We believe that many of these adaptations correspond to optimal solutions driven by evolution, training, and underlying physical and chemical laws and constraints. While some argue against such optimality principles due to their potential ambiguity, we propose generalized inverse optimal control to infer them directly from data. This novel approach incorporates multi-criteria optimality, nestedness of objective functions on different scales, the presence of active constraints, the possibility of switches of optimality principles during the observed time horizon, maximization of robustness, and minimization of time as important special cases, as well as uncertainties involved with the mathematical modeling of biological systems. This data-driven approach ensures that optimality principles are not merely theoretical constructs but are firmly rooted in experimental observations. Furthermore, the inferred principles can be used in forward optimal control to predict and manipulate biological systems, with possible applications in bio-medicine, biotechnology, and agriculture. As discussed and illustrated, the well-posed problem formulation and the inference are challenging and require a substantial interdisciplinary effort in the development of theory and robust numerical methods.
@@ 1 Introduction
We shall generalize this concept within the realm of optimization by interpreting “model” as an optimization model $\min\phi(y)$ subject to $y\in\mathcal{Y}$. With this interpretation, we can then make forward predictions for observations given $\phi$ and $\mathcal{Y}$. Conversely, the inverse optimization problem infers the unknown components of $\phi$ and $\mathcal{Y}$ from given observations. The forward car driving scenario might involve determining a policy for accelerating, switching gears, and steering a car efficiently navigate from $A$ to $B$ in minimal time, while adhering to constraints imposed by the track and engine speed. Conversely, the inverse optimization problem entail observing the car’s position and velocity and determining whether the driver is prioritizing energy-minimal or time-minimal driving.
@@
Our main goal is to provide an overview of the use of optimality principles in the biological sciences, focusing specifically on the cellular level. We discuss how these principles have been and could be used to explain various biological processes and systems. We highlight the main challenge unique to this field: unlike in previously mentioned contexts, the functions being optimized in biology are often unknown a priori. Simultaneously, it is highly desirable to have at least approximate knowledge of these functions wether for exploitation in an engineering context or for scientific insight. To address this issue and tackle several criticism often raised against the general consideration of optimality in biology within the literature, we present the novel generalized inverse optimal control framework. This framework builds upon previous approaches of inverse optimal control, extending them towards a more general inference of complete optimization models. These models involve constraints, partially unknown dynamics, and possible changes in modus operandi.
This paper follows the following structure: Section 2 delves into optimality principles in biology, including a historic survey of concepts and related discussions. Readers can approach this section with the intuitive notion of optimality as something that represents the best among all alternatives, while inverse optimality is framed as the task to inferring optimality from data. In Section 3 we introduce the novel methodology of generalized inverse optimal control, providing a more technical exploration. Here, we also offer an introductory survey of important concepts and definitions in optimization and optimal control. Finally, the paper concludes with Section 4, where we summarize our findings and provide concluding remarks.
@@ 2 Optimality principles in biology
Biology encompasses a vast array of scales, ranging from the molecular level to planetary ecosystems. On one hand, one might be tempted to extend principles such as least action or maximization of entropy from physics and chemistry to all these biological scales. On the other hand, emergence is a well-known phenomenon, dating back to Anderson’s seminal paper More is different from 1972 [13]. Presently, the emergence across scales is widely acknowledged [14]. Therefore, a more nuanced examination of optimality in biology is warranted. This is particularly crucial given the processes of training and the evident evolutionary advantage associated with optimality, which serve as additional contributing factors.
@@ 2.1.2 Optimal control
Optimality in the vascular system [104, 105, 106]
@@
Optimality in sensorimotor control [120, 121]
@@ 2.3.2 Uses
In other words, optimization theory in biology allow us to assess our knowledge and understanding of the evolutionary trajectory and diversity of life forms, the mechanisms within evolved biological systems, and how they will respond to new conditions [126, 162, 102]. For instance, one key insight from the comparisons of bacterial genomes is that their “lifestyle” (as an adaptation to their environment) significantly influences their genomes [166]. A notable example is that long-term mutualists (endosymbionts) of insects, living in very stable environments, have the smallest genomes [166, 171]. Studying these minimal natural genomes can be helpful to estimate the least amount of genetic components needed to construct a contemporary, free-living cellular entity, a key step to create a living cell [171], the ultimate goal of synthetic biology.
@@ 2.3.3 Terminology
Previous discussions and disputes regarding the application of optimality principles in biology may have stemmed from confusion surrounding various optimization terminologies, or the employment ambiguously defined terms. For instance, the precise definitions and boundaries of concepts like “constraints”, “trade-offs”, “robustness”, and “evolvability” have not always been consistently established, contributing to the lack of clarity in some discussions around optimality principles. In this context, it can be helpful to discuss and clarify in more detail several key concepts, starting with causality, constraints, and trade-offs. Mayr introduced the concepts of proximate and ultimate causation in biology [172]. Proximate causation refers to the immediate mechanisms underlying a biological trait (e.g., resource limitations). Ultimate causation refers to the evolutionary processes that shape a trait (e.g., ecological circumstances). Proximate causes operate within an organism’s lifetime, while ultimate causes involve Darwinian selection across generations. These two concepts can help us understand trade-offs in biology [173]: ultimate trade-offs can act through proximate mechanisms and those mechanisms can evolve over time. In other words, understanding both the how and the why of a trait can provide a more complete picture of its evolution. Further, as discussed by Birch [163], it is important to distinguish optimization principles that concern what happens at equilibrium (e.g., explaining a biological system after a long period of evolution in a constant environment) from those that concern the direction of change in evolving species.
@@
Such trade-offs can be addressed using mathematical multi-criteria optimization, where Pareto optimality is a key concept. It describes a situation where the value of one objective function cannot be improved without impairing at least one other objective function. The Pareto optimal set represents the set of solutions where there are inherent trade-offs between conflicting objectives. Therefore, multi-criteria optimization captures the best compromises between conflicting objectives and constraints, providing a systematic and quantitative framework for identifying the set of optimal solutions that balance these trade-offs. Pareto optimality has been successfully used in various areas of biology, including evolutionary, systems, and synthetic biology, see [39, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183] and references therein for further details.
@@ 2.4 Key assumptions
These and related criticisms motivate the development of new methods aimed at inferring optimality principles directly from data. While the call for a reverse optimality approach was made several decades ago [20, 126], a comprehensive methodology of this nature is still lacking. We believe that the time is ripe for the development of data-driven approaches capable of analyzing extensive biological dynamic datasets to uncover the fundamental principles dictating these dynamics, while integrating constraints from the outset. In other words, we claim the necessity for a method for unbiased identification of optimality principles from dynamic (time-series) data and prior knowledge. This approach would ensure that these optimality principles are not merely speculative but grounded in empirical evidence and quantifiable constraints.
To ensure that such a method addresses most of the critiques outlined above, it would be beneficial to consider several assumptions and clarifications.
Firstly, we assume that the optimality principle can be mathematically represented as a constrained multicriteria dynamic optimization problem. This formulation yields Pareto optimal solutions, providing a natural means of managing the interplay between constraints and trade-offs. The selection of appropriate elementary objective functions is problem-dependent and should prioritize those with strong selection pressure, defined as the functions expected to have the most significant impact on fitness. We are also assuming that dynamics play a key role in the optimality principle. While network structure and steady state data can define key aspects of function, robustness, and regulation in certain cases (e.g., metabolic pathways [202]), it is generally challenging to determine a priori if this alone is sufficient to infer optimality. Therefore, we posit that a kinetic mechanistic description is necessary.
Second, active constraints can significantly influence the optimization process: if a constraint is binding (i.e., it is satisfied as an equality at the optimal solution), it can create a trade-off. Relaxing the constraint may enable improvement in one objective without compromising another, thereby altering the trade-off. In many cases, these constraints are unknown a priori and must be inferred from the available data.
@@ 3 Generalized Inverse Optimal Control
For a better intuition of the properties of the problem class, we shall discuss its components one by one and examine special cases that are relevant in biology in Section 3.1. We will illustrate the concept by providing examples in Section 3.2. Finally, we discuss numerical methods for solving (1) and review previous work in IOC and IRL in Section 3.3.
@@ 3.1.1 Dynamic systems and variables $y$
We start with a closer look at the variables $y$ in (1). In mathematical modeling, we usually distinguish between state (dependent) and control (independent) variables. State variables, denoted here by $x$, represent the internal state (past, current, and future condition) of the system. Control variables, denoted by $u$ and also known as input variables, have slightly different meanings depending on the application. In engineering and design, controls are the variables that can be manipulated or controlled to achieve a desired output or response from the system. In the natural sciences, they are the external forces or signals that affect the system’s behavior (i.e., as described by its states). The mathematical model links $x$ and $u$, allowing to calculate the dependent variables $x$ for any choice of $u$. There are many different ways to model a dynamic system, all of which can be combined with problem (1), such as Markov chains, physics-informed neural networks [205], or partial stochastic delay differential-algebraic equations. Particularly relevant for molecular systems biology are ordinary differential equations, which we will focus on in the following for a clearer representation, but without loss of generality. Therefore, the general variable $y$ in (1) will be specified to variables $(x,u)$. Often, the mechanics of a dynamic system can be captured with an initial value problem of the form
@@
and that allow to modify values such as reaction rates, masses, initial concentrations or sizes, coefficients of Michaelis-Menten kinetics, or similar. The underlying assumption is, though, that a priori biological knowledge (how do states $x$ and controls $u$ relate to one another independent of the specific values of $p$) can be included in the function $f(\cdot)$. If this is not the case, there are various options for learning the relationships from data. For example, parts of the function $f(\cdot)$ may consist of neural networks (referred to as universal differential equation [208]) or algebraic functions can be learned via symbolic regression [204]. Here, the function $f$ (or parts of it) can be an arbitrary mathematical function composed from basis functions such as multiplication, exponential, and addition of $x$ and $u$. Although different algorithmic approaches are necessary as discussed in Section 3.3, all of these different modeling approaches can be conveniently represented via formulation (5) with a general function $f(\cdot)$ depending on a parameter vector $p$. From a modeling perspective, we shall thus not further discuss the nuances of numerical regression, hybrid systems, and symbolic regression.
@@ 3.1.2 The objective function $\phi$
For particular applications, it is interesting to consider multiple objectives $\phi_{i}$. Often, there is a trade-off, e.g., between time and energy consumption to achieve a certain goal. A point $y^{*}\in\mathcal{Y}$ is said to be not dominated if there is no other $y\in\mathcal{Y}$ exists such that $\phi_{i}(y)\leq\phi_{i}(y^{*})$ for all $i$, i.e., which is better with respect to all objectives. The set of all non-dominated points is called the Pareto front. It is of particular interest for a posteriori decision-making, allowing the balancing advantages and disadvantages of the different objective functions $\phi_{i}$.
@@
Among all possible objective function candidates, two stand out as particularly relevant but also as special cases concerning modeling. The first one is the minimization of time. From a mathematical point of view, treating $T$ as an optimization variable (formally, this can be easily achieved via a model parameter $p_{i}$ multiplied to the right-hand side of the ODE) implies another, problem-dependent modeling choice: how shall observed values $\eta_{i}$ at times $t>T$ be treated? One can either discard them, compare them to $h(y^{*}(T))$, simulate $y$ for times $t>T$ with some assumption on the applied controls, or force $T$ to be smaller than the maximum observation time. The second one is the maximization of robustness. As discussed above, robustness refers to the ability of a biological system to maintain its function despite perturbations [184, 185]. In robust optimization and optimal control many different approaches have been proposed, e.g., [210, 211, 212]. Important concepts include those of random and targeted attacks. Two detailed case studies of how parametric uncertainty can be used in biological networks are described in [213]. The authors compare linearization, sigma points, and polynomial chaos approaches. In general, all of them can be used within our gIOC approach by extending the mathematical model with additional states. These correspond either to variational differential equations (i.e., sensitivities of the states with respect to parameters) or to sample points, allowing an estimation of higher order moments of the underlying distributions. These higher-order moments can then be used to model candidate functions $\phi_{i}$ or candidate constraint functions $g_{i}$ that represent aspects of robustness. In summary, robustness can be considered as a special case of function candidates, albeit at the expense of additional modeling effort and higher computational costs to solve problem (8). If objective function candidates are not known, this can again be addressed via formulations including neural networks or via symbolic regression. In analogy to (5), this case is formally included via a dependence of the functions $\phi_{i}$ on the model parameter vector $p$. Wether all candidates are known or not is application-dependent. While the general formulation (1) covers both scenarios, different algorithms and special care considering issues of identifiability are necessary.

### v2

@@ Abstract
Living organisms exhibit remarkable adaptations across all scales, from molecules to ecosystems. We believe that many of these adaptations correspond to optimal solutions driven by evolution, training, and underlying physical and chemical laws and constraints. While some argue against such optimality principles due to their potential ambiguity, we propose generalized inverse optimal control to infer them directly from data. This comprehensive approach incorporates multi-criteria optimality, nestedness of objective functions on different scales, the presence of active constraints, the possibility of switches of optimality principles during the observed time horizon, maximization of robustness and minimization of time as important special cases, as well as uncertainties involved with the mathematical modeling of biological systems. This data-driven approach ensures that optimality principles are not merely theoretical constructs but are firmly rooted in experimental observations. The inferred principles can also be used in forward optimal control to predict and manipulate biological systems, with possible applications in bio-medicine, biotechnology, and agriculture. As discussed and illustrated, the well-posed problem formulation and the inference are challenging and require a substantial interdisciplinary effort in the development of theory and robust numerical methods.
@@ 1 Introduction
We shall generalize this concept within the realm of optimization by interpreting “model” as an optimization model $\min\phi(y)$ subject to $y\in\mathcal{Y}$. With this interpretation, we can then make forward predictions for observations given $\phi$ and $\mathcal{Y}$. Conversely, the inverse optimization problem infers the unknown components of $\phi$ and $\mathcal{Y}$ from given observations. The forward car driving scenario might involve determining a policy for accelerating, switching gears, and steering a car efficiently to navigate from $A$ to $B$ in minimal time, while adhering to constraints imposed by the track and engine speed. Conversely, the inverse optimization problem entails observing the car’s position and velocity and determining whether the driver is prioritizing energy-minimal or time-minimal driving.
@@
Our main goal is to provide an overview of the use of optimality principles in the biological sciences, focusing specifically on the cellular level. We discuss how these principles have been and could be used to explain various biological processes and systems. We highlight the main challenge unique to this field: unlike in previously mentioned contexts, the functions being optimized in biology are often unknown a priori. Simultaneously, it is highly desirable to have at least approximate knowledge of these functions whether for exploitation in an engineering context or for scientific insight. To address this issue and tackle several criticism often raised against the general consideration of optimality in biology within the literature, we present the generalized inverse optimal control framework. This framework builds upon previous approaches of inverse optimal control, extending them towards a more general inference of complete optimization models. These models involve constraints, partially unknown dynamics, and possible changes in modus operandi.
This paper has the following structure: Section 2 delves into optimality principles in biology, including a historic survey of concepts and related discussions. Readers can approach this section with the intuitive notion of optimality as something that represents the best among all alternatives, while inverse optimality is framed as the task to inferring optimality from data. In Section 3 we introduce the methodology of generalized inverse optimal control, providing a more technical exploration. Here, we also offer an introductory survey of important concepts and definitions in optimization, optimal control, and inverse reinforcement learning. Finally, the paper concludes with Section 4, where we summarize our findings and provide concluding remarks.
@@ 2 Optimality principles in biology
Biology encompasses a vast array of scales, ranging from the molecular level to planetary ecosystems. On the one hand, one might be tempted to extend principles such as least action or maximization of entropy from physics and chemistry to all these biological scales. On the other hand, emergence is a well-known phenomenon, dating back to Anderson’s seminal paper More is different from 1972 [13]. Presently, the emergence across scales is widely acknowledged [14]. Therefore, a more nuanced examination of optimality in biology is warranted. This is particularly crucial given the processes of training and the evident evolutionary advantage associated with optimality, which serve as additional contributing factors.
@@ 2.1.2 Optimal control
Optimality in the vascular system [106, 107, 108, 109]
@@
Optimality in sensorimotor control [123, 124, 125, 126]
@@ 2.3.2 Uses
In other words, optimization theory in biology allows us to assess our knowledge and understanding of the evolutionary trajectory and diversity of life forms, the mechanisms within evolved biological systems, and how they will respond to new conditions [131, 166, 104]. For instance, one key insight from the comparisons of bacterial genomes is that their “lifestyle” (as an adaptation to their environment) significantly influences their genomes [170]. A notable example is that long-term mutualists (endosymbionts) of insects, living in very stable environments, have the smallest genomes [170, 175]. Studying these minimal natural genomes can be helpful to estimate the least amount of genetic components needed to construct a contemporary, free-living cellular entity, a key step to create a living cell [175], the ultimate goal of synthetic biology.
@@ 2.3.3 Terminology
Previous discussions and disputes regarding the application of optimality principles in biology may have stemmed from confusion surrounding various optimization terminologies, or the employment of ambiguously defined terms. For instance, the precise definitions and boundaries of concepts like “constraints”, “trade-offs”, “robustness”, and “evolvability” have not always been consistently established, contributing to the lack of clarity in some discussions around optimality principles. In this context, it can be helpful to discuss and clarify in more detail several key concepts, starting with causality, constraints, and trade-offs. Mayr introduced the concepts of proximate and ultimate causation in biology [176]. Proximate causation refers to the immediate mechanisms underlying a biological trait (e.g., resource limitations). Ultimate causation refers to the evolutionary processes that shape a trait (e.g., ecological circumstances). Proximate causes operate within an organism’s lifetime, while ultimate causes involve Darwinian selection across generations. These two concepts can help us understand trade-offs in biology [177]: ultimate trade-offs can act through proximate mechanisms and those mechanisms can evolve over time. In other words, understanding both the how and the why of a trait can provide a more complete picture of its evolution. Further, as discussed by Birch [167], it is important to distinguish optimization principles that concern what happens at equilibrium (e.g., explaining a biological system after a long period of evolution in a constant environment) from those that concern the direction of change in evolving species.
@@
Such trade-offs can be addressed using mathematical multi-criteria optimization, where Pareto optimality is a key concept. It describes a situation where the value of one objective function cannot be improved without impairing at least one other objective function. The Pareto optimal set represents the set of solutions where there are inherent trade-offs between conflicting objectives. Therefore, multi-criteria optimization captures the best compromises between conflicting objectives and constraints, providing a systematic and quantitative framework for identifying the set of optimal solutions that balance these trade-offs. Pareto optimality has been successfully used in various areas of biology, including evolutionary, systems, and synthetic biology, see [39, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188] and references therein for further details.
@@ 2.4 Key assumptions
These and related criticisms motivate the development of new methods aimed at inferring optimality principles directly from data. While the call for a reverse optimality approach was made several decades ago [20, 131], a comprehensive methodology of this nature is still lacking. We believe that the time is ripe for the development of data-driven approaches capable of analyzing extensive biological dynamic datasets to uncover the fundamental principles dictating these dynamics, while integrating constraints from the outset. In other words, we claim the necessity for a method for unbiased identification of optimality principles from dynamic (time-series) data and prior knowledge. This approach would ensure that these optimality principles are not merely speculative but grounded in empirical evidence and quantifiable constraints. To ensure that such a method addresses most of the critiques outlined above, it would be beneficial to consider several assumptions and clarifications.
First, we assume that the optimality principle can be mathematically represented as a constrained multicriteria dynamic optimization problem. This formulation yields Pareto optimal solutions, providing a natural means of managing the interplay between constraints and trade-offs. The selection of appropriate elementary objective functions is problem-dependent and should prioritize those with strong selection pressure, defined as the functions expected to have the most significant impact on fitness. We are also assuming that dynamics play a key role in the optimality principle. While network structure and steady state data can define key aspects of function, robustness, and regulation in certain cases (e.g., metabolic pathways [207]), it is generally challenging to determine a priori if this alone is sufficient to infer optimality. Therefore, we posit that a kinetic mechanistic description is necessary.
Second, active constraints can significantly influence the optimization process: if a constraint is binding (i.e., it is satisfied as an equality at the optimal solution), it can create a trade-off. Relaxing the constraint may enable improvement in one objective without compromising another, thereby altering the trade-off. Often these constraints are unknown a priori and must be inferred from the available data.
@@ 3 Generalized Inverse Optimal Control
For a better intuition of the properties of the problem class, we shall discuss its components one by one and examine special cases that are relevant in biology in Section 3.1. We will illustrate the concept by providing examples in Section 3.2. We discuss numerical methods for solving (1) and review the state-of-the-art in IOC and IRL in Section 3.3. We provide links to valuable previous work especially in the robotics and model predictive control communities, which may serve as a good basis for future gIOC methods for applications in biology.
@@ 3.1.1 Dynamic systems and variables $y$
We start with a closer look at the variables $y$ in (1). In mathematical modeling, we usually distinguish between state (dependent) and control (independent) variables. State variables, denoted here by $x$, represent the internal state (past, current, and future condition) of the system. Control variables, denoted by $u$ and also known as input variables, have slightly different meanings depending on the application. In engineering and design, controls are the variables that can be manipulated or controlled to achieve a desired output or response from the system. In the natural sciences, they are the external forces or signals that affect the system’s behavior (i.e., as described by its states). The mathematical model links $x$ and $u$, allowing to calculate the dependent variables $x$ for any choice of $u$. There are many different ways to model a dynamic system, all of which can be combined with problem (1), such as Markov chains, physics-informed neural networks [210], or partial stochastic delay differential-algebraic equations. Particularly relevant for molecular systems biology are ordinary differential equations (ODE), which we will focus on in the following for a clearer representation, but without loss of generality. Therefore, the general variable $y$ in (1) will be specified to variables $(x,u)$. Often, the mechanics of a dynamic system can be captured with an initial value problem of the form
@@
and that allow to modify values such as reaction rates, masses, coefficients of Michaelis-Menten kinetics, initial concentrations or sizes, or similar. The underlying assumption is, though, that a priori biological knowledge (how do states $x$ and controls $u$ relate to one another independent of the specific values of $p$) can be included in the function $f(\cdot)$. If this is not the case, there are various options for learning the relationships from data. For example, parts of the function $f(\cdot)$ may consist of neural networks (referred to as universal differential equation [213]) or algebraic functions can be learned via symbolic regression [209]. Here, the function $f$ (or parts of it) can be an arbitrary mathematical function composed from basis functions such as multiplication, exponential, and addition of $x$ and $u$. Although different algorithmic approaches are necessary as discussed in Section 3.3, all of these different modeling approaches can be conveniently represented via formulation (5) with a general function $f(\cdot)$ depending on a parameter vector $p$. From a modeling perspective, we shall thus not further discuss the nuances of numerical regression, hybrid systems, and symbolic regression.
@@ 3.1.2 The objective function $\phi$
For particular applications, it is interesting to consider multiple objectives $\phi_{i}$. Often, there is a trade-off, e.g., between time and energy consumption to achieve a certain goal. A point $y^{*}\in\mathcal{Y}$ is said to be not dominated if there is no other $y\in\mathcal{Y}$ exists such that $\phi_{i}(y)\leq\phi_{i}(y^{*})$ for all $i$, i.e., which is better with respect to all objectives. The set of all non-dominated points is called the Pareto front. It is of particular interest for a posteriori decision-making as it allows balancing the advantages and disadvantages of the different objective functions $\phi_{i}$.
@@
Among all possible objective function candidates, two stand out as particularly relevant but also as special cases concerning modeling. The first one is the minimization of time. From a mathematical point of view, treating $T$ as an optimization variable (formally, this can be easily achieved via a model parameter $p_{i}$ multiplied to the right-hand side of the ODE) implies another, problem-dependent modeling choice: how shall observed values $\eta_{i}$ at times $t>T$ be treated? One can either discard them, compare them to $h(y^{*}(T))$, simulate $y$ for times $t>T$ with some assumption on the applied controls, or force $T$ to be smaller than the maximum observation time. The second one is the maximization of robustness. As discussed above, robustness refers to the ability of a biological system to maintain its function despite perturbations [189, 190]. In robust optimization and optimal control different approaches have been proposed, e.g., [215, 216, 217]. Important concepts include those of random and targeted attacks. Two detailed case studies of how parametric uncertainty can be used in biological networks are described in [218]. The authors compare linearization, sigma points, and polynomial chaos approaches. In general, all of them can be used within our gIOC approach by extending the mathematical model with additional states. These correspond either to variational differential equations (i.e., sensitivities of the states with respect to parameters) or to sample points, allowing an estimation of higher order moments of the underlying distributions. These higher-order moments can then be used to model candidate functions $\phi_{i}$ or candidate constraint functions $g_{i}$ that represent aspects of robustness. In summary, robustness can be considered as a special case of function candidates, albeit at the expense of additional modeling effort and higher computational costs to solve problem (8). If objective function candidates are not known, this can again be addressed via formulations including neural networks or via symbolic regression. In analogy to (5), this case is formally included via a dependence of the functions $\phi_{i}$ on the model parameter vector $p$. Whether all candidates are known or not is application-dependent. While the general formulation (1) covers both scenarios, different algorithms and special care considering issues of identifiability are necessary.